# Zero-Shot Tokenizer Transfer

### v1

@@ Appendix A Unigramifying: Approximating Arbitrary Tokenizers via UnigramLM
$\mathcal{L_{T}}(X,\hat{T})=\sum_{x\in X}\,\sum_{C\in\mathcal{C}_{x}\setminus\{T(x)\}}\max\left(0,\sum_{t\in T(x)}\log p_{\hat{T}}(t)-\sum_{t\in C}\log p_{\hat{T}}(t)\right)$
@@ Appendix C Non-Amortizing Hypernetworks
We experimented with hypernetworks taking the tokenization function into account by adding sparse inter-token attention blocks between the self-attention and the FFN in every hypernetwork layer. Sparse inter-token attention consists of two attention blocks. The first attention block attends from a fixed amount of learnable inter-token embeddings (e.g. 16, each a vector of size $d_{\text{model}}$) to the $i$th token representation of every token sequence passed to the hypernetwork. The second block attends from the $i$th token representation to the inter-token embeddings. This way, we factorize the attention to e.g. one $16\times k$ attention and one $k\times 16$ attention, instead of regular the $k\times k$ self-attention which would be infeasibly slow for typical vocabulary sizes. We only add inter-token attention for the first token in every sequence. This improves performance on the sampled tokenizers, but does not improve performance on ‘real-world’ tokenizers (Table 9); investigating this mismatch is a direction for future work.
@@ Appendix D Additional Hyperparameters
Hyperparameters for hypernetwork training are shown in Table 10. For continued training, we use the same optimizer, but sequence length of 512, batch size of 32, training for 50k steps and a constant learning rate chosen among the set $\{1\mathrm{e}{-6},3\mathrm{e}{-6},6\mathrm{e}{-6},1\mathrm{e}{-5},3\mathrm{e}{-5}\}$ to maximize performance. The chosen learning rate is $1\mathrm{e}{-6}$ for the runs keeping the original tokenizer (original@800M), $6\mathrm{e}{-6}$ for continued training starting from FOCUS (FOCUS@800M) and $3\mathrm{e}{-6}$ for continued training with the hypernetwork (ours@800M).
@@ Appendix F Reliance on Vocabulary Overlap
Appendix G Additional LLM Results
@@
Appendix H Assumptions on the Tokenization Function

### v2

@@ 6 Conclusion
7 Limitations
The key limitation of our approach is the requirement to train a hypernetwork for every base model. Although the hypernetwork only needs to be trained once, doing so is computationally intensive and may not be feasible for many LLM practitioners. Instead, it may be a task LLM providers are better positioned to undertake. Other limitations are the remaining assumptions on the tokenization function (Appendix I), and not taking the tokenization function into account (Appendix J), although these limitations do not appear to have substantial impact in practice. Finally, we have limited our scope to experiments on text-only models, but Zero-Shot Tokenizer Transfer could also be beneficial for multimodal models, such as models ‘perceiving’ images or speech; we leave this to future work.
@@ Appendix A Unigramifying: Approximating Arbitrary Tokenizers via UnigramLM
$\mathcal{L_{T}}(X,\hat{T})=\sum_{x\in X}\,\sum_{C\in\mathcal{C}_{x}\setminus\{T(x)\}}\max\left(0,\sum_{t\in C}\log p_{\hat{T}}(t)-\sum_{t\in T(x)}\log p_{\hat{T}}(t)\right)$
@@ Appendix C Non-Amortizing Hypernetworks
We experimented with hypernetworks taking the tokenization function into account by adding sparse inter-token attention blocks between the self-attention and the FFN in every hypernetwork layer. Sparse inter-token attention consists of two attention blocks. The first attention block attends from a fixed amount of learnable inter-token embeddings (e.g. 16, each a vector of size $d_{\text{model}}$) to the $i$th token representation of every token sequence passed to the hypernetwork. The second block attends from the $i$th token representation to the inter-token embeddings. This way, we factorize the attention to e.g. one $16\times k$ attention and one $k\times 16$ attention, instead of the standard $k\times k$ self-attention which would be infeasibly slow for typical vocabulary sizes. We only add inter-token attention for the first token in every sequence. This improves performance on the sampled tokenizers, but does not improve performance on ‘real-world’ tokenizers (Table 8); investigating this mismatch is a direction for future work.
@@ Appendix D Additional Hyperparameters
Hyperparameters for hypernetwork training are shown in Table 9. For continued training, we use the same optimizer, but a sequence length of 512, batch size of 32, training for 50k steps and a constant learning rate chosen among the set $\{1\mathrm{e}{-6},3\mathrm{e}{-6},6\mathrm{e}{-6},1\mathrm{e}{-5},3\mathrm{e}{-5}\}$ to maximize performance. The chosen learning rate is $1\mathrm{e}{-6}$ for the runs keeping the original tokenizer (original@800M), $6\mathrm{e}{-6}$ for continued training starting from FOCUS (FOCUS@800M) and $3\mathrm{e}{-6}$ for continued training with the hypernetwork (ours@800M).
@@ Appendix F Reliance on Vocabulary Overlap
Appendix G Reliance on Overlap between Hypernet Training Tokens and Target Tokens
Figure 6: Analyzing how often the hypernetwork sees the tokens of different target tokenizers during training. Note the logarithmic y-scale. We analyze the occurrence for all tokens in the target vocabulary (top) and for tokens which occur at least once in the evaluation data (bottom) across target tokenizers in seen languages for XLM-R (left), unseen XLM-R languages (middle) and English Mistral-7B tokenizers (right). The bottom row is more informative w.r.t. how well the hypernetwork generalizes to unseen tokens since tokens which do not occur do not substantially impact evaluation.
@@ Appendix G Reliance on Overlap between Hypernet Training Tokens and Target Tokens
We analyze how often the hypernetwork sees the tokens in the vocabulary of different target tokenizers across multiple settings in Figure 6. We differentiate between tokens which occur in the evaluation data, and tokens which do not; this is important since the embeddings of tokens which do not occur in the evaluation data will not substantially impact performance. Notably, for XLM-R, >35% of occurring tokens in Greek, Bulgarian and Russian are unseen by the hypernet, even though the hypernet is trained on these languages. This is likely due to the non-Latin scripts. The hypernet still performs well in these languages with an average 2% performance decrease at 17% sequence length reduction on XNLI. In total, the HN has seen approx. 200M different tokens during training.
Appendix H Additional LLM Results
@@
Appendix I Assumptions on the Tokenization Function
@@
Appendix J Tokenization Function Amortization and Unigramifying Results
Results measuring the success of unigramifying tokenizers are shown in Table 12. Results measuring the success of amortizing over the tokenization function are shown in Table 13.
BERT Mistral-7B TinyLlama-1.1B GPT2
Kind WordPiece BPE BPE BBPE
Original $p(\text{preserved})$ 100% 100% 100% 100%
bits per char n/a 0.675 0.747 0.930
To Byte-Level $p(\text{preserved})$ 99.6% 99.9% 99.9% 100%
Extra Tokens 162 522 362 0
Unigramify $p(\text{preserved})$ 99.4% 99.8% 99.8% 99.7%
bits per char n/a 0.678 0.750 0.932
Model Embeddings Tokenizer $\bm{(\mathcal{V},T)}$
$(\text{GPT2},\text{GPT2})$ $(\text{GPT2},\text{unigramify}(\text{GPT2}))$ $(\text{GPT2},\text{UnigramLM})$
GPT2 original 0.930 0.932 1.005
ours 0.919 0.920 0.964
ours (no noise) 0.925 0.926 0.978
Model Hypernet
#params FLOPs / token #params FLOPs / token
GPT2 124M 253M 21M (16%) 4.5M (1.8%)
TinyLlama-1.1B 1.1B 2.1G 170M (15%) 33.1M (1.6%)
Mistral-7B 7.2G 15.4G 678M   (9%) 132.1M (0.9%)
Appendix K Analyzing FLOPs of the hypernetwork
Estimated FLOPs per token for the hypernet and the corresponding main model are shown in Table 14. We estimate FLOPs on the basis of XLA-compiled instructions using Jax (Bradbury et al., 2018).